QUESTION 7

or th below two machines and based on CC analysis which machine we should select? MARR=10%.


Machine AMachine B
First cost, $              24,311          100,000
Annual cost, $/year                 13,553               7,000
Salvage value, $                 6,323 - 
Life, years 3infinite


Answer the below question:

A- the CC for machine A= 
Answer #1

SOLUTION :


r = MARR = 10% = 0.1

=> (1 + r) = 1.1


PV of costs of Machine A :


= 24311 + 13353/ 1.1 + 13353/1.1^2 + 13353/1.1^3 - 6323/1.1^3 

= 52767.37 ($) 


PV of costs of Machine B :


= 100000 + 7/0.1

= 100070 ($) 


Therefore, as per CC, Machine A has lower cost. So Machine A should be selected.

(ANSWER)

answered by: Tulsiram Garg

> Please correct : 7th line as : = 100000+ 7000/0.1 ; 8th line as : = 170000 ($)

Tulsiram Garg Mon, Oct 11, 2021 2:32 AM